# Hardware and Software Platform Inference

### v1

@@ Abstract
It is now a common business practice to buy access to large language model (LLM) inference rather than self-host, because of significant upfront hardware infrastructure and energy costs. However, as a buyer, there is no mechanism to verify the authenticity of the advertised service including the serving hardware platform, e.g. that it is actually being served using an NVIDIA H100. Furthermore, there are reports suggesting that model providers may deliver models that differ slightly from the advertised ones, often to make them run on less expensive hardware. That way, a client pays premium for a capable model access on more expensive hardware, yet ends up being served by a (potentially less capable) cheaper model on cheaper hardware. In this paper we introduce hardware and software platform inference (HSPI) – a method for identifying the underlying GPU architecture and software stack of a (black-box) machine learning model solely based on its input-output behavior. Our method leverages the inherent differences of various GPU architectures and compilers to distinguish between different GPU types and software stacks. By analyzing the numerical patterns in the model’s outputs, we propose a classification framework capable of accurately identifying the GPU used for model inference as well as the underlying software configuration. Our findings demonstrate the feasibility of inferring GPU type from black-box models. We evaluate HSPI against models served on different real hardware and find that in a white-box setting we can distinguish between different GPUs with between $83.9\%$ and $100\%$ accuracy. Even in a black-box setting we are able to achieve results that are up to three times higher than random guess accuracy.
@@ 1 Introduction
To this end, we introduce two methods: HSPI with Border Inputs (HSPI-BI) and HSPI with Logits Distributions (HSPI-LD). We demonstrate the efficacy of these techniques in HSPI under both white-box and black-box setups, extending across both vision and language tasks. Our work demonstrates the practicality of HSPI and examines its performance for various models running over diverse device families, showcasing the influence of system optimizations and variations within software and hardware supply chains.
Overall, we make the following contributions:
  • We define the Hardware and Software Platform Inference (HSPI), detailing the underlying assumptions.
@@ 2 What makes HSPI possible?
In this section we discuss why it is possible to infer the hardware and software configuration of a machine learning serving platform just from input-output pairs. While familiar to those in high-performance computing, we explain how varying software and hardware configurations can shift a model into different Equivalence Classes, and how arithmetic ordering and optimizations can contribute to computational discrepancies.
Equivalence Classes: Different hardware and software configurations give us various computational results. When the computational results stay the same and do not deviate between settings, we talk about them being in the same equivalence class (EQC). EQCs are used to group similar computational behaviors that yield consistent results under specific settings, such as quantization levels, GPU architectures, CUDA versions, and batch sizes. For instance, quantization can alter numerical precision and thereby shift a model into a different EQC, resulting in subtle variations in outputs. GPU architecture and batch size also affect precision consistency, as different hardware or parallelizing data processing methods introduce minor deviations in results. Schlögl et al. provide a more granular analysis, examining EQCs layer by layer across models, identifying how architectural choices impact computational stability and precision deviations (Schlögl et al., 2023).
Factors Influencing Computational Deviations: Schlögl et al. describe several possible reasons for precision deviations which include arithmetic units using faster approximations, intermediate values being rounded to fit in registers of different sizes, transformations made during execution planning, and various tricks that optimize performance in ML toolboxes such as loop unrolling, constant folding, and arithmetic simplifications (Schlögl et al., 2021). In a follow-up paper they show that the order of aggregation and choice of convolution algorithm can largely affect computation results (Schlögl et al., 2023). A simple example for aggregation errors leading to varying results when we are rounding all decimals to integers is given by following equations:
@@
Similar examples can be found in any quantization due to finite numerical precision and stochastic pipelining.
Furthermore, various methods exist to optimize the calculation of convolutional layers. Convolution calculation can for example be generalized to matrix multiplications. The space of calculations can also be transformed to, e.g. the Fourier domain, which then allows convolutions to be replaced by multiplication. Varying implementations for convolutions include the direct loop, GEMM (NVIDIA, 2024), Winograd (Winograd, 1980) and Fourier transform (Smith, 1997).
Schlögl et al. find that the extent of deviations in computations is influenced by several factors within neural network architectures. Layers involving high multiplication counts per convolution tend to amplify deviations, as they accumulate rounding errors and precision loss. Additionally, data-preserving layer types, which retain the full spectrum of input data, contribute to increased deviations, while layers that aggregate data (e.g., pooling layers) help mitigate these errors. Precision choice also plays a critical role: single-precision operations show a higher likelihood of deviation than both double and half precision due to their reduced bit representation (Schlögl et al., 2023).
@@ 3 Related Work
This section reviews relevant prior literature and positions our work within it.
@@
Hardware Fingerprinting in ML: Schlögl et al. first identify unique hardware fingerprints, discovering that different hardware platforms are in different EQCs (Schlögl et al., 2021, 2021). This lays the groundwork to our hardware fingerprinting methods for GPUs. They discuss boundary samples, which are inputs to a model that output results at the decision boundary between two classes. In classification for example an input can lie at the boundary of being classified as cat or dog, depending on the hardware.
They use an adaptation of the search algorithm iterative fast-gradient-sign-method (FGSM) also known as Projected Gradient Descent (PGD) for generating adversarial samples and differentiate between 4 CPU models. They develop their method to be half remote, as they argue that it is more realistic to assume that only one CPU micro-architecture can be accessed locally and instantly, while more latency would be involved in communicating with others. This is why boundary sample generation is split into a local and remote phase, where the local phase only involves finding an input very close to a boundary and the remote phase checks whether this sample flips the classification label only on one microarchitecture. However, this algorithm is inefficient due to the two phase setup and the remote phase has a success rate of creating a border sample of only $28.25\%$ on CIFAR10. Furthermore, GPUs and a black-box setting are not considered, i.e., it is assumed that the ML model that is being run is known. Additionally, the analyses contain only models in the visual domain.
Model Inference Variability: Further investigations into convolutional neural networks (CNNs) by Schlögl et al. reveal that numerical deviations in calculations arise from variations in SIMD usage and convolution algorithms (Schlögl et al., 2023). Their analysis serves as a foundation for our research, providing insight into how these deviations can be exploited in model inference time.
Model Training Variability: Additionally, related research also shows big variances in model training in terms of accuracy and training time. Even after excluding variations introduced through random seeds, training configurations, deep learning libraries, and hardware factors, an accuracy difference of up to $2.9\%$ can be introduced through software implementation differences of deep learning (Pham et al., 2020). Others formulate this problem as a repeatability problem and underline the non-determinism in the whole training stack greatly impacting model performance and safety (Zhuang et al., 2022; Alahmari et al., 2020).
@@
Exploitation of Floating Point Inaccuracies: Some other work has also emerged exploiting floating point inaccuracies, for example, against robustness verifiers of neural networks (Jia and Rinard, 2021; Zombori et al., 2021). They point out that errors in floating point can be used against verifiers that do not consider deviations in floating point computations. While Zombori et al. suggest adding small perturbations to the weights as an adhoc mitigation, they warn against other attacks using these deviations. As such, Clifford et al. for instance show that deviations in how specific operations are calculated on different hardware platforms can enable locking models to certain hardware (Clifford et al., 2024).
@@
Black-box LLM Model Identification: Other works have focused on identifying the LLM model family from a black-box setting with only access to output text through semantic analysis (Iourovitski et al., 2024; Pasquini et al., 2024; McGovern et al., 2024), a variant where logit outputs are assumed (Yang and Wu, 2024) or a more white-box setting where outputs from each layer can be accessed (Zeng et al., 2023; Zhang et al., 2024). Accuracies for distinguishing between model families in the black-box setting have been $72\%$ and $95\%$ with as little as 8 interactions (Iourovitski et al., 2024; Pasquini et al., 2024) which shows that we can create a more white-box setting to start our LLM hardware platform inference with little probing.
@@ 4 Methodology
This section begins by outlining HSPI formally and listing the underlying assumptions. Subsequently, we introduce the concept of border inputs, which are inputs that are specifically designed to elicit divergent behavior across varying hardware and software configurations. Finally, we describe a method for model differentiation that is achievable even with not-specifically-crafted data, by identifying deviations in floating-point distribution of model-returned logits.
@@ 4.1 Problem Formulation and Assumptions
The implementation of model serving in practical settings differs in accessibility. Service providers like Google, OpenAI, or Anthropic keep the underlying model architecture undisclosed, and users simply send queries and receive responses through API calls. Most services, including OpenAI APIs, are capable of providing users with the model output probabilities (logits). In other scenarios, we have cloud providers, such as Azure and AWS, offering deployment of open-source models, like LLaMA (Dubey et al., 2024), where both the model weights and architectures are known. We thus evaluate the following two representative scenarios:
  • White-box access to the model – the deployed model is known and can be accessed locally. For example, the deployed model is available publicly e.g. LLaMA (Dubey et al., 2024) or Gemma (Team et al., 2024). It is thus possible in an unrestricted way to send inputs and receive responses from the model locally;
  • Black-box access only – the deployed model can only be accessed via the input interface, where it returns the output probabilities. This is similar to the current serving practices of Google, OpenAI, or Anthropic.
We also assume that it is possible to access $N$ different hardware platforms (=$\mathcal{H}=\{H_{0},H_{1}...,H_{N-1}\}$), where HSPI then tries to identify the hardware $H_{t}\in\mathcal{H}$ that the model is deployed on. To the best of our knowledge, currently, all known hardware platforms suitable for model serving can be rented, making our approach feasible and realistic.
As newer hardware devices get released, such as recent Groq (Gwennap, 2020) and Cerebras (Lie, 2022), they quickly become privately rentable through either cloud providers or directly from respective hardware vendors. We expect the list of platforms $\mathcal{H}$ to grow and newer devices to eventually become accessible on-demand.
It is important to note that slight performance and numerical variations may arise from the underlying software stack, even though the algorithm and hardware remain constant. For example, in context of machine learning GPU kernel fusion strategies and runtime scheduling can influence the EQC of the model. We also explicitly include these variations in our $\mathcal{H}$. Consequently, HSPI can also be used to determine both the hardware platform and the software configuration, thereby identifying the combined hardware-software supply chain.
@@ 4.2 HSPI-BI: HSPI with Border Inputs
We reintroduce the concept of boundary samples with the name border inputs. As explained by Schlögl et al. these are specially crafted inputs that are at the decision boundary between two or more output classes of a model (Schlögl et al., 2023). Similarly to Schlögl et al. we find that the idea of border inputs is similar to the idea of adversarial examples and we also modify PGD to create border inputs but formulate our loss function differently.
Specifically, consider a model $F$ which runs in two different hardware environments. These environments could for example differ in hardware device, quantization level, or batch size. We recognize that deploying models on alternative hardware environments $H$ or $H^{\prime}$ might lead to subtle divergences in logit outputs $F_{H}$ due to discrepancies in quantization level, the sequence of GPU arithmetic operations (Schlögl et al., 2023), or disparities in the core library implementations (eg. SDPA vs. FlashAttention). To find a point of maximum logit discrepancy we can start changing an input $X$ such that the two logit outputs $F_{H}(X)$ and $F_{H^{\prime}}(X)$ become as different as possible, with the goal that ultimately the predicted class labels $y$ and $y^{\prime}$ differ. We can construct such a loss by defining it as the sum of two cross-entropy loss terms, each comparing the output logits of one model with the target labels from the other model. Specifically, we define the loss $L_{pgd}$ as the cross entropy loss $\mathcal{L}$ between the output logits of $F_{H}$ and the predicted class label $y^{\prime}$ of $F_{H^{\prime}}$, and the cross entropy loss between the output logits of $F_{H^{\prime}}$ and predicted class label $y$ of $F_{H}$, i.e.,
$L_{pgd}=\mathcal{L}(F_{H}(X),y^{\prime})+\mathcal{L}(F_{H^{\prime}}(X),y).$ (3)
The loss should grow as the difference between $F_{H}(X)$ and $F_{H^{\prime}}(X)$ increases. To encourage the models to predict different class labels, $y$ and $y^{\prime}$, we aim to maximize this loss. For this, we adjust the input by moving it slightly in the direction that maximizes the loss, as determined by the gradient. At each step, we also clamp the modified input within a valid range to keep it close to the original input.
Another approach to increasing the distance between $F_{H}(X)$ and $F_{H^{\prime}}(X)$ is to push $F_{H}$ closer to a specific target class, while pushing $F_{H^{\prime}}$ further away from that same target class. We can achieve this by defining the loss as the difference between the cross entropy loss for $F_{H}(X)$ with respect to the target label $y_{t}$, and the cross-entropy loss for $F_{H^{\prime}}(X)$ with the same target label. By adjusting the input in the direction of the negative gradient, the loss is minimized:
@@
To generate inputs that are classified differently across various hardware environments, we can extend Equation 4. For all models on different hardware setups, we sum the cross-entropy loss between the target label and each model’s logits. Then, for the one model where we want a distinct class, we subtract the loss with the target label:
$L_{\text{1-vs-rest, targeted}}=\left(\sum_{k\in\mathcal{N}\setminus\{i\}}({\mathcal{L}(F_{H_{i}}(X),y_{t})}\right)-\mathcal{L}(F_{H}(X),y_{t})).$ (5)
As the number of hardware environments increases, optimizing between models becomes more challenging. In such cases, we can apply Equation (3) iteratively across pairs of models to estimate the most likely hardware environment probabilistically. Furthermore, when we consider testing various border inputs in a black box setting we do not necessarily need access to logits anymore. In order to determine the hardware environment probabilistically with a set of border inputs only the output class is needed.
4.3 HSPI-LD: HSPI with Logits Distributions
An alternative approach to executing the HSPI task involves developing a classification model by using the distribution of output logits. The idea is that the logits of a set of inputs should reveal characteristics of the hardware environment. The logit distribution is especially revealing if inputs are used that are very diverse in distribution of classes and closeness to class boundaries. To further emphasize these differences, we can convert the logits into binary representations, or more cost-efficiently, split the floating-point components (Figure 2). The theory is that models running on different hardware configurations may produce distinct bit patterns in their logits, with certain bits used more or less frequently depending on the hardware. Given access to all hardware configurations $\mathcal{H}$ and input samples $X_{i}\in\mathcal{X}$, we can create a classifier $G$ that learns to identify the environment (e.g., whether a sample was processed by $F_{H}$ or $F_{H^{\prime}}$). Mathematically, this classifier can be defined as:
@@ 5 Evaluation
In this section, we first describe the experiment setup in Section 5.1, and then present the results of white-box and black-box attacks in Section 5.2 and Section 5.3 respectively.
@@ 5.1 Experiment Setup
We consider the white-box and black-box attack setups described in Section 4.1. We build the set of hardware $\mathcal{H}$ under two distinct configurations: (1) Initially, we examine models in low-precision formats (quantization), a popular technique integrated into widely-used platforms like HuggingFace (Wolf et al., 2020), TensorRT (NVIDIA, 2024), and TorchAO (Torchao, 2024). This is because hardware devices typically feature specific arithmetic operators (e.g., INT4 for A100 and FP8 for H100), and we regard differentiating different low-precision formats as a preliminary step before differentiating actual GPUs. (2) We then extend the experiments to a setup close to real deployment scenarios, comparing actual GPUs to operate with various arithmetic configurations, different kernel implementations, and across varying device families.
@@
We mainly use the image classification models from torchvision and fine-tune them on CIFAR10, including ResNet18, ResNet50 (He et al., 2016), VGG16 (Simonyan and Zisserman, 2014), EfficientNet-B0 (Tan and Le, 2019), DenseNet-121 (Fu et al., 2021), MobileNet-v2 (Sandler et al., 2018), MobileNet-v3-small, and MobileNet-V3-large (Howard et al., 2019). Detailed fine-tuning parameters and the fine-tuned accuracy on the test split of CIFAR10 can be found in Section A.1.
We perform the experiments on standard floating-point formats (FP32, FP16, BF16) and low-precision formats first, then extend to actual GPUs. The low-precision formats include MXINT8 (Darvish Rouhani et al., 2020), FP8-E3M4 (noted as FP8-E3) (Shen et al., 2024), FP8-E4M3 (noted as FP8-E4), and dynamic INT8 (noted as INT8) (Kim et al., 2021). We show in Section A.2 that all low-precision formats have negligible accuracy loss compared to the original model checkpoints. For actual GPUs, we include NVIDIA H100, A100, GeForce RTX2080 Ti, and Quadro RTX8000. Not all GPUs where used for all experiments due to different server locations and the difficulty of connecting all of them for the attack generation phase.
@@
We use the open-sourced instruction-tuned LLMs from HuggingFace, including DistillGPT2 (Sanh et al., 2019), LLaMA-3.1-8B (Dubey et al., 2024), QWen-2.5-1.5B/3B/7B (Yang et al., 2024), Phi-3.5-mini (Abdin et al., 2024), Gemma-2-2B/9B (Team et al., 2024), and Mistral-7B-v0.3 (Jiang et al., 2023). Using the official chatting templates, these instruction-tuned models can be directly deployed to chatbot applications.
Similar to vision models, we first run experiments of differentiating low-precision formats, then extend to actual GPUs. Since LLM quantization is more challenging than vision models, besides FP16 and BF16, we adopt two quantization methods that has been proven effective and integrated into HuggingFace: fine-grained dynamic INT8 quantization (Torchao, 2024), noted as INT8-FD, and half-quadratic quantization (Badri and Shaji, 2023), noted as HQQ-4bit. For actual GPUs, we consider NVIDIA A100, L40S, RTX A6000, and GeForce RTX3090.
For HSPI-BI method, we use random texts as the initial border requests, and apply the PGD method to update the border requests. A border request can be represented as a list of vocabulary indexes (input IDs). Since input IDs are one-hot encoded, we need to ensure that the updated border requests are still valid input IDs at the start of each PGD step. To achieve this, we first update the one-hot encoded input IDs, then use argmax to find the index of the maximum value in the updated vector for each token. For HSPI-LD method, we send artificial requests to guide the model to generate random words. Figure 3 shows an example of the border requests and the generated responses. We append several random words to each request. The logits of generated tokens (response) plus the quantization level/GPU form a pair of input and label to train an SVM machine. Since the vocabulary size varies across model families, we sample the generated tokens and logits.

### v2

@@ Abstract
It is now a common business practice to buy access to large language model (LLM) inference rather than self-host, because of significant upfront hardware infrastructure and energy costs. However, as a buyer, there is no mechanism to verify the authenticity of the advertised service including the serving hardware platform, e.g. that it is actually being served using an NVIDIA H100. Furthermore, there are reports suggesting that model providers may deliver models that differ slightly from the advertised ones, often to make them run on less expensive hardware. That way, a client pays premium for a capable model access on more expensive hardware, yet ends up being served by a (potentially less capable) cheaper model on cheaper hardware. In this paper we introduce hardware and software platform inference (HSPI) – a method for identifying the underlying GPU architecture and software stack of a (black-box) machine learning model solely based on its input-output behavior. Our method leverages the inherent differences of various GPU architectures and compilers to distinguish between different GPU types and software stacks. We evaluate HSPI against models served on different real hardware and find that in a white-box setting we can distinguish between different GPUs with between $83.9\%$ and $100\%$ accuracy. Even in a black-box setting we achieve results that are up to 3$\times$ higher than random guess accuracy. Our code is available at https://github.com/ChengZhang-98/HSPI.
Security, LLM, GPU Architecture, Model Serving
@@ 1 Introduction
To this end, we introduce two methods: HSPI with Border Inputs (HSPI-BI) and HSPI with Logits Distributions (HSPI-LD). We demonstrate the efficacy of these techniques in HSPI under both white-box and black-box setups, extending across both vision and language tasks. Overall, we make the following contributions:
  • We define Hardware and Software Platform Inference (HSPI), detailing the underlying assumptions.
@@ 2 What makes HSPI possible?
We briefly explain how varying software and hardware configurations can shift a model into different Equivalence Classes, and how arithmetic ordering and optimizations can contribute to computational discrepancies.
Equivalence Classes: Different hardware and software configurations give us various computational results. When the computational results stay the same and do not deviate between settings, we talk about them being in the same equivalence class (EQC). EQCs are used to group similar computational behaviors that yield consistent results under specific settings, such as quantization levels, GPU architectures, and batch sizes. Schlögl et al. (2023) provide a more granular analysis, examining EQCs layer by layer across models, identifying how architectural choices impact computational stability and precision deviations.
[…]
@@
Due to finite numerical precision and stochastic pipelining, such examples can be found at any quantization level. Quantization is a technique for improving hardware efficiency at inference time. Additionally, Schlögl et al. (2023) find that the extent of deviations in computations is influenced by neural network architectures, e.g., layers involving high multiplication counts tend to amplify deviations.
@@ 3 Related Work
Hardware Fingerprinting in ML: Schlögl et al. first identify unique hardware fingerprints, discovering that different hardware platforms are in different EQCs (2021; 2021). They propose boundary samples, which are inputs to a model that output results at the decision boundary between two classes. For example, an input can lie at the boundary of being classified as cat or dog, depending on the hardware. They use an adaptation of the search algorithm iterative fast-gradient-sign-method (FGSM), also known as Projected Gradient Descent (PGD), for generating adversarial samples and differentiate between 4 CPU models.
This sets the groundwork for our hardware fingerprinting methods. However, their algorithm is inefficient due to the two phase setup and the remote phase has a success rate of only $28.25\%$ on CIFAR10. Our work improves this method with HSPI-BI, and we additionally propose a new method without the need of backward propagation (HSPI-LD). Further, we mainly target GPUs, considering both white-box setup and black-box setup, and extensively perform evaluations on modern ML workloads like large language models.
@@
Exploitation of Floating Point Inaccuracies: Some work has also emerged exploiting floating point inaccuracies, for example, against robustness verifiers of neural networks (Jia & Rinard, 2021; Zombori et al., 2021). They point out that errors in floating point can be used against verifiers that do not consider deviations in floating point computations. Zombori et al. suggest adding small perturbations to the weights as an adhoc mitigation, they warn against other attacks using these deviations. As such, Clifford et al. for instance show that deviations in how specific operations are calculated on different hardware platforms can enable locking models to certain hardware (Clifford et al., 2024).
@@
Black-box LLM Model Identification: Other works have focused on identifying the LLM model family in a black-box setting with only access to output text through semantic analysis (Iourovitski et al., 2024; Pasquini et al., 2024; McGovern et al., 2024), a variant where logit outputs are assumed (Yang & Wu, 2024) or a more white-box setting where outputs from each layer can be accessed (Zeng et al., 2023; Zhang et al., 2024). Accuracies for distinguishing between model families in the black-box setting have been $72\%$ and $95\%$ with only 8 interactions (Iourovitski et al., 2024; Pasquini et al., 2024).
@@ 4 Methodology
We start by formally describing the HSPI and listing the underlying assumptions in Section 4.1. In Section 4.2, we propose border inputs, which are inputs that are specifically designed to elicit divergent behavior in varying hardware and software configurations. In Section 4.3, we devise another method without the need of specifically-crafted data, by identifying deviations in floating-point distribution of model-returned logits.
@@ 4.1 Problem Formulation and Assumptions
The implementation of model serving in practical settings differs in accessibility. Service providers like Google, OpenAI, or Anthropic keep the underlying model architecture undisclosed, and users simply send queries and receive responses through API calls. Most services, including OpenAI APIs, are capable of providing users with the model output probabilities (logits). In other scenarios, we have cloud providers, such as Azure and AWS, for the deployment of open-source models, where both the model weights and architectures are known. We thus evaluate the following two representative scenarios:
  • White-box access to the model – the deployed model is known and can be accessed freely. For example, publicly available LLMs, e.g., LLaMA (Dubey et al., 2024) and Gemma (Team et al., 2024), can be deployed locally for sending queries and receiving responses;
  • Black-box access only – the deployed model can only be accessed via cloud-based interface, where it returns the output probabilities. This is similar to the current serving practices of Google, OpenAI, and Anthropic.
We also assume that it is possible to access $N$ different hardware platforms (=$\mathcal{H}=\{H_{0},H_{1}...,H_{N-1}\}$), where HSPI then tries to identify the hardware $H_{t}\in\mathcal{H}$ that the model is deployed on. To the best of our knowledge, currently, all known hardware platforms suitable for model serving can be rented. Even newer hardware devices eventually are becoming accessible on demand as seen with recent Groq (Gwennap, 2020) and Cerebras (Lie, 2022) hardware, thus making our approach feasible and realistic.
Note that slight performance and numerical variations may arise from the underlying software stack, even though the algorithm and hardware remain constant. For example, in context of machine learning, GPU kernel fusion strategies and runtime scheduling can influence the EQC of the model. We also explicitly include these variations in our $\mathcal{H}$. Consequently, HSPI can also be used to determine both the hardware platform and the software configuration, thereby identifying the combined hardware-software supply chain.
@@ 4.2 HSPI-BI: HSPI with Border Inputs
We reintroduce the concept of boundary samples with the name border inputs. As explained by Schlögl et al. these are specially crafted inputs that are at the decision boundary between two or more output classes of a model (2023). The idea of border inputs is similar to the idea of adversarial examples and we also modify PGD to create border inputs, however, formulate our loss function differently.
Specifically, consider a model $F$ which runs in two different hardware environments. Deploying models on alternative hardware environments $H$ or $H^{\prime}$ can lead to subtle divergences in logit outputs $F_{H}$. To maximize this divergence, we modify input $X$ until the predicted labels $y$ and $y^{\prime}$ differ. We define a loss function $L_{pgd}$ as the sum of cross-entropy losses between each model’s logits and the other’s predicted labels:
$\mathcal{L}_{pgd}=\mathcal{L}(F_{H}(X),y^{\prime})+\mathcal{L}(F_{H^{\prime}}(X),y).$ (3)
We maximize $L_{pgd}$ by iteratively updating $X$ along the gradient direction while clamping it within a valid range of the original input, thereby encouraging the models to predict different class labels.
We can also maximize divergence by pushing $F_{H}(X)$ toward a target class $y_{t}$ while pushing $F_{H^{\prime}}(X)$ away from it. This is achieved through a loss function that subtracts the cross-entropy losses:
@@
When the number of possible serving platforms $N$ is small, we can extend this to compare one model against all others by summing all losses and subtracting the loss only for the one model for which we want a distinct class:
$\mathcal{L}_{\text{1-vs-rest}}=\left(\sum_{k\in\mathcal{N}\setminus\{i\}}({\mathcal{L}(F_{H_{i}}(X),y_{t})}\right)-\mathcal{L}(F_{H}(X),y_{t})).$ (5)
As $N$ increases, optimizing between models becomes more challenging. In such cases, we decompose the HSPI task as $\frac{N(N-1)}{2}$ binary classification problems by applying Equation 3 or Equation 4 iteratively across pairs of models to estimate the most likely hardware environment probabilistically. Note that when testing various precomputed border inputs in a black-box setting, logits become unnecessary and only the predicted class is needed.
4.3 HSPI-LD: HSPI with Logit Distributions
An alternative approach involves developing a classification model leveraging the information in the distribution of output logits. The logits of a set of inputs reveal characteristics of the hardware environment. The logit distribution is especially informative when inputs are diverse in the distribution of classes and closeness to class boundaries. To amplify these characteristics, we convert the logits into binary representations for small models, or more cost-efficiently, split the floating-point components (Figure 2) for large models.
We find models running on different hardware configurations produce distinct bit patterns in their logits, with certain bits used more or less frequently. Given access to all hardware configurations $\mathcal{H}$ and input samples $X_{i}\in\mathcal{X}$, we can create a classifier $G$ that learns to identify the environment (e.g., whether a sample was processed by $F_{H}$ or $F_{H^{\prime}}$). Mathematically, this classifier can be defined as:
@@ 4.3 HSPI-LD: HSPI with Logit Distributions
4.4 Model-Specific Adaptation
We adapt the general methods defined in Section 4.2 and Section 4.3 to fit vision models or language models. For vision models, the border inputs of HSPI-BI are randomly sampled from the model’s training set, while the input images of HSPI-LD are random noise as they trigger more diverse areas of the feature space. We also carefully round border images or noise images to integers between 0 and 255 to ensure the final images are available in PNG format.
For language models, we generate random texts as the initial border queries of HSPI-BI. Since input IDs are one-hot encoded, we need to ensure that the updated border requests are still valid input IDs at the start of each PGD step. To achieve this, we first update the one-hot encoded input IDs, then use argmax to find the index of the maximum value in the updated vector for each token. For the HSPI-LD method, we guide the model to generate random words and sample top-$k$ logits of each token. We find such prompts encourage the top-$k$ logits to have closer values and avoid -Inf values. Figure 3 shows how output logits are sampled.
@@ 5 Evaluation
In this section, we briefly describe the experiment setup in Section 5.1, and then present the results of white-box in Section 5.2 and black-box attacks  Section 5.3.
@@ 5.1 Experiment Setup
We consider the white-box and black-box attack setups described in Section 4.1. We build the set of hardware $\mathcal{H}$ under two distinct configurations: (1) Initially, we examine models in low-precision formats (quantization), a popular technique integrated into frameworks like HuggingFace (Wolf et al., 2020), TensorRT (NVIDIA, 2024), and TorchAO (Torchao, 2024). Hardware devices typically feature specific arithmetic operators (e.g., INT4 for A100 and FP8 for H100), and we regard differentiating different low-precision formats as a preliminary step before differentiating actual GPUs. (2) We then extend the experiments to a setup closer to real deployment scenarios, comparing actual GPUs to operate with various arithmetic configurations, different kernel implementations, and across varying device families.
@@
We mainly use the image classification models from torchvision and fine-tune them on CIFAR10, including ResNet18, ResNet50 (He et al., 2016), VGG16 (Simonyan & Zisserman, 2014), EfficientNet-B0 (Tan & Le, 2019), DenseNet-121 (Fu et al., 2021), MobileNet-v2 (Sandler et al., 2018), MobileNet-v3-small, and MobileNet-V3-large (Howard et al., 2019). We shortlist the number formats, and GPU specs in Section B.1. Note that not all GPUs were used for all groups of experiments due to different server locations and the difficulty of connecting all of them for the attack generation phase.
@@
We use the open-sourced instruction-tuned LLMs from HuggingFace, including DistillGPT2 (Sanh et al., 2019), LLaMA-3.1 (Dubey et al., 2024), QWen-2.5 (Yang et al., 2024), Phi-3.5 (Abdin et al., 2024), Gemma-2 (Team et al., 2024), and Mistral (Jiang et al., 2023). We use the official chat templates to guide the LLM to generate random words. The prompts also include a few manually added random words to increase the diversity of generation. Detailed prompts can be found in Section B.2. Besides standard floating-point formats, we adopt two SoTA quantization methods highly optimized for LLMs. The detailed configurations and GPU specs can also be found in Section B.2.